# On the stability of a penalty function of the $\mathbb{Z}^{2}$-hard square shift

### v1

@@ Abstract.
We investigate the stability of maximizing measures for a penalty function of a two-dimensional subshift of finite type, building on the work of Gonschorowski et al. [GQS]. In the one-dimensional case, such measures remain stable under Lipschitz perturbations for any subshift of finite type. However, instability arises in a penalty function of the Robinson tiling, which is a two-dimensional subshift of finite type with no periodic point and zero entropy. This raises the question of whether stability persists in two-dimensional subshifts of finite type with positive topological entropy. In this paper, we address this question by studying the $\mathbb{Z}^{2}$-hard square shift, a well-known example of a two-dimensional subshift with positive entropy. Our main theorem establishes that, in contrast to previous results, a penalty function of the hard square shift remains stable under Lipschitz perturbations.
@@ 1. Introduction
where $\mathcal{M}_{T}(X)$ is the space of $T$-invariant Borel probability measures on $X$ endowed with the weak*-topology. An invariant measure which attains the maximum is called a maximizing measures for $\varphi$ and denote by $\mathcal{M}_{{\rm max}}(\varphi)$ the set of maximizing measures for $\varphi$.
@@ 2. Settings
for $\underline{x},\underline{y}\in\mathcal{A}^{\mathbb{Z}^{2}}$ where $i=\inf\{\|u\|_{\infty}:x_{u}\neq y_{u}\}$. Then, this metric is compatible with the given topology.
@@ 3. Proof of the main theorem
Set $I=f^{-1}\{0\}$. Since the set of maximizing measures is convex and closed, it suffices to prove the statement for ergodic measures. Let $\mu$ be an ergodic invariant measure supported on $X^{c}$.
@@
which complete the proof.
@@
Let $\underline{x}$ be a generic point of $\mu$. For each $i\in\mathbb{Z}$ let
@@
$\tilde{x}_{(u_{1},u_{2})}=\left\{\begin{array}[]{cc}0&\mbox{if}\ u_{1}\in S^{k},u_{2}=k\ \mbox{and}\ k\in\mathbb{Z}\\ x_{(u_{1},u_{2})}&\mbox{othewise}.\end{array}\right.$
Clearly, $\tilde{x}\in X$. By definition we have $\lim_{k\to\infty}\underline{x}^{(k)}=\tilde{x}$.

### v2

# On the stability of the penalty function for the $\mathbb{Z}^{2}$-hard square shift

@@ Abstract.
We investigate the stability of maximizing measures for a penalty function of a two-dimensional subshift of finite type, building on the work of Gonschorowski et al. [GQS]. In the one-dimensional case, such measures remain stable under Lipschitz perturbations for any subshift of finite type. However, instability arises for a penalty function of the Robinson tiling, which is a two-dimensional subshift of finite type with no periodic point and zero entropy. This raises the question of whether stability persists in two-dimensional subshifts of finite type with positive topological entropy. In this paper, we address this question by studying the $\mathbb{Z}^{2}$-hard square shift, a well-known example of a two-dimensional subshift with positive entropy. Our main theorem establishes that, in contrast to previous results, a penalty function of the hard square shift remains stable under Lipschitz perturbations.
@@ 1. Introduction
where $\mathcal{M}_{T}(X)$ is the space of $T$-invariant Borel probability measures on $X$ endowed with the weak*-topology. An invariant measure which attains the maximum is called a maximizing measure for $\varphi$ and denote by $\mathcal{M}_{{\rm max}}(\varphi)$ the set of maximizing measures for $\varphi$.
@@
We remark that the stability result for the hard square shift is relatively straightforward, since forbidden words can be easily eliminated by replacing $1$ with 00. However, extending this result to other nearest neighbor subshifts of finite type, such as those discussed in [Pav12], appears to be more challenging. This difficulty arises because, in general, there is no method to identify the exact locations of forbidden words, making it impossible to estimate the distances required to remove them.
@@ 2. Settings
for $\underline{x},\underline{y}\in\mathcal{A}^{\mathbb{Z}^{2}}$ where $i=\inf\{\|u\|_{\infty}:x_{u}\neq y_{u}\}$. Then, this metric is compatible with the product topology.
@@ 3. Proof of the main theorem
Set $I=f^{-1}\{0\}$. Since the set of maximizing measures is convex and closed, it suffices to prove the result for ergodic measures. Let $\mu$ be an ergodic invariant measure supported on $X^{c}$.
@@
which completes the proof.
@@
Let $\underline{x}$ be a generic point for the measure $\mu$. For each $i\in\mathbb{Z}$ let
@@
$\tilde{x}_{(u_{1},u_{2})}=\left\{\begin{array}[]{cc}0&\mbox{if}\ u_{1}\in S^{k},u_{2}=k\ \mbox{and}\ k\in\mathbb{Z}\\ x_{(u_{1},u_{2})}&\mbox{otherwise}.\end{array}\right.$
Clearly, $\tilde{x}\in X$. By definition, we have $\lim_{k\to\infty}\underline{x}^{(k)}=\tilde{x}$.